OpenStudy (anonymous):

Use the logarithm rules to combine into single log expression: 2lnx + 3lny - lnz

OpenStudy (bahrom7893):

lnA + lnB = Ln(AB)

OpenStudy (bahrom7893):

lnA-lnB = Ln(A/B)

OpenStudy (bahrom7893):

aLnb = lnb^a

myininaya (myininaya):

is that a duck?

OpenStudy (bahrom7893):

lol no it's a boomerang bird

myininaya (myininaya):

oh lol

OpenStudy (bahrom7893):

So according to those ln rules, u have: lnx^2 + lny^3 - lnz

OpenStudy (bahrom7893):

So: ln(x^2y^3/z)

myininaya (myininaya):

oops lol

myininaya (myininaya):

gj bahrom!

OpenStudy (bahrom7893):

U TOO MYININAYA!